# Inference for multiple treatment effects using confounder importance learning

### v4

@@ H.1. Illustration of the EB and EP objective functions
Figure 9. Empirical Bayes (left) and Expectation-Propagation (right) objective functions (13) and (14) in the single treatment case ($T=1$). Here, ${\bm{\theta}}^{\textsc{eb}}=(-2.43,3.19)$ and ${\bm{\theta}}^{\textsc{ep}}=(-2.34,3.09)$, for $n=100$ and $J=49$, for the first data realization for the simulation design displayed in the center-left panel of Figure 1 with three confounders. See Section 4.1 for further details.
Figure 9 shows the Empirical Bayes objective function in (13) and (14) in a simulated dataset with a single treatment. A bimodality is appreciated in the left panel.
H.2. Salary survey: obtention and pre-processing of CPS microdata
@@
H.3. Salary survey: generation of augmented datasets
@@
H.4. Further results on salary survey
@@
H.5. Further results for abortion data
@@
We first discuss violent crime. Tables H.5 and 5 show BMA inference for covariates with marginal posterior inclusion probability above $>0.1$ in the standard BMA analysis with normal and MOM priors (respectively). In the CIL analyses with normal and MOM priors there are no such covariates. The top model in the BMA-normal, BMA-MOM, CIL-normal and CIL-MOM analysis contained no confounding covariates, and has posterior probabilities of 0.095, 0.305, 0.923 and 0.670 respectively.
Regarding property crime, the middle panels in Tables H.5 and 5 show results for BMA-Normal and BMA-MOM respectively. Tables 6 and 7 show analogous results for CIL-Normal and CIL-MOM.
Finally, for the murder outcome the results from our CIL methodology are very similar to those from standard BMA. The bottom panels in Tables H.5 and 5 show results for murder under the BMA-Normal and BMA-MOM analyses. The covariate with highest posterior inclusion probability is a term related to the quadratic effect of income. Said covariate is also the only one receiving non-negligible posterior inclusion probability under the CIL analyses (Tables 6-7). In fact, the top model under all analyses contained only this covariate and has a posterior probability of 0.465 for BMA-normal, 0.612 for BMA-MOM, 0.437 for CIL-normal and 0.802 for CIL-MOM.
@@
[…]
H.6. Model selection results in simulation study
@@
Figure 11 summarizes model selection results for the simulations described in Figure 1.
H.7. Simulations under growing dimensionality ($T=1$)
@@
H.8. Testing CIL to different amounts of confounders for $T=1$
@@
Figure 13 shows the effect of having various amounts of active confounders. The results look consistent to the effects reported in Figures 1 and 12, which are magnified for large amounts of active confounders. These are really challenging situations to tackle since the tested methods aim at model sparsity, while the true model size is relatively large. Although our method still performed at oracle rates in low-confounding scenarios, its relative performance is compromised for the highest levels of confounding. This occurred in part because accurate point estimation in (5) become increasingly harder as the correlation between covariates strengthened, which in turn influenced the ability of the algorithm to calibrate ${\bm{\theta}}$ reliably. Even in these hard cases, however, its performance is not excessively far to the best competing method, while it clearly outperformed BMA on all of them.

### v6

@@ 15.1. Illustration of the EB and EP objective functions
Figure 9. Empirical Bayes (left) and Expectation-Propagation (right) objective functions (3.8) and (3.9) in the single treatment case ($T=1$). Here, ${\bm{\theta}}^{\textsc{eb}}=(-2.43,3.19)$ and ${\bm{\theta}}^{\textsc{ep}}=(-2.34,3.09)$, for $n=100$ and $J=49$, for the first data realization for the simulation design displayed in the center-left panel of Figure 1 with three confounders. See Section 5.1 for further details.
Figure 9 shows the Empirical Bayes objective function in (3.8) and (3.9) in a simulated dataset with a single treatment. A bimodality is appreciated in the left panel.
15.2. Salary survey: obtention and pre-processing of CPS microdata
@@
15.3. Salary survey: generation of augmented datasets
@@
15.4. Further results on salary survey
@@
Table 2 provides a descriptive analysis of the covariates in the salary data that changed the most between 2010 and 2019. These are covariates with a $p$-value$<0.05$ when assessing their marginal correlation with year (based on a linear for non-binary outcomes, and a chi-squared test for binary outcomes). Further, we only report covariates whose average changed by at least 5% (in absolute value) between 2010 and 2019. For binary covariates, we also required that their average in 2010 was $>0.05$, to discard covariates that were very rare (or not collected) in 2010. The variable codes used match those supplied by the CPS database.
Covariate Label Type Prop. change Mean (2010)
(2010-19)
Change of industry from last year chindly binary $0.093$ $0.105$
Change of occupation from last year choccly binary $0.051$ $0.066$
Has pension plan at work pension binary $-0.130$ $0.507$
Employment coverage (family plan) grptyply_family binary $-0.177$ $0.362$
Employment coverage (self plan) grptyply_self binary $-0.141$ $0.247$
Medical out-of-pocket & Medicare B subsidy (log) logspmmedxpns non-binary $0.105$ $6.645$
Mortage mortgage binary $-0.101$ $0.588$
Size of firm where employed (25-99 people) firmsize_2(25-99) binary $-0.055$ $0.131$
Lunch subsidy lunchsub binary $0.052$ $0.078$
Metropolitan area size $>5$ areasize_6(>5) binary $0.063$ $0.159$
Proportion of income from interest propoi_incint non-binary $0.529$ $0.295$
Proportion of income from dividends propoi_incdivid non-binary $-0.423$ $0.075$
Weeks unemployed last year (log) wksunemly non-binary $-0.510$ $1.296$
Amount of child tax credit (log \$) logctccrd non-binary $0.265$ $1.566$
Number of children with school lunch subsidy frelunch non-binary $0.570$ $0.142$
Log of other person’s income logotherpersincome non-binary $0.117$ $3.326$
Number of children who ate school lunch atelunch non-binary $-0.125$ $0.523$
Proportion of tax income over wages (log) logproptaxincwageinc non-binary $-0.062$ $1.124$
Number of children (log) logspmnchild non-binary $-0.073$ $0.972$
Number of own children nchild non-binary $-0.062$ $1.083$
Number of own children under age 5 nchlt5 non-binary $-0.107$ $0.215$
Family market value of school lunch (log) logschllunch non-binary $-0.068$ $1.426$
Number of siblings nsibs non-binary $0.193$ $0.080$
Number of months receiving food stamps stampmo non-binary $0.147$ $0.463$
Table 2. Descritive analysis for covariates that changed the most in the salary data between 2010 and 2019. These all displayed a $p$-value$<0.05$ for their marginal association with year, and changed at least by 5% in absolute value between 2010 and 2019. Label indicates the name of the variable in our processed dataset
15.5. Further results for abortion data
@@
We first discuss violent crime. Tables LABEL:tab:abortion_bma_normal and LABEL:tab:abortion_bma_mom show BMA inference for covariates with marginal posterior inclusion probability above $>0.1$ in the standard BMA analysis with normal and MOM priors (respectively). In the CIL analyses with normal and MOM priors there are no such covariates. The top model in the BMA-normal, BMA-MOM, CIL-normal and CIL-MOM analysis contained no confounding covariates, and has posterior probabilities of 0.095, 0.305, 0.923 and 0.670 respectively.
Regarding property crime, the middle panels in Tables LABEL:tab:abortion_bma_normal and LABEL:tab:abortion_bma_mom show results for BMA-Normal and BMA-MOM respectively. Tables 7 and 8 show analogous results for CIL-Normal and CIL-MOM.
Finally, for the murder outcome the results from our CIL methodology are very similar to those from standard BMA. The bottom panels in Tables LABEL:tab:abortion_bma_normal and LABEL:tab:abortion_bma_mom show results for murder under the BMA-Normal and BMA-MOM analyses. The covariate with highest posterior inclusion probability is a term related to the quadratic effect of income. Said covariate is also the only one receiving non-negligible posterior inclusion probability under the CIL analyses (Tables 7-8). In fact, the top model under all analyses contained only this covariate and has a posterior probability of 0.465 for BMA-normal, 0.612 for BMA-MOM, 0.437 for CIL-normal and 0.802 for CIL-MOM.
@@
Dinc2 $-91217.96$ $-604869.53$ 0.00 0.18
15.6. Further results for the simulation study
Bias2
$\alpha=1$ $\alpha=1/3$ $\alpha=0$
Refer to caption Refer to caption Refer to caption
Variance
$\alpha=1$ $\alpha=1/3$ $\alpha=0$
Refer to caption Refer to caption Refer to caption
Figure 11. Squared bias and variance for the simulation scenarios described in Figure 1 considering strong ($\alpha=1$), weak $(\alpha=1/3)$ and no effect $(\alpha=0)$. The x-axis quantifies the amount of confounding, measured by the number of covariates that are truly related to both the outcome and the treatment
$\alpha=1$ $\alpha=1/3$ $\alpha=0$
Refer to caption Refer to caption Refer to caption
Figure 12. Sensitivity to tuning parameter in BAC and ACPME. Parameter root MSE relative to an oracle OLS for the simulation scenarios described in Figure 1 considering strong ($\alpha=1$), weak $(\alpha=1/3)$ and no effect $(\alpha=0)$
$\alpha=1$ $\alpha=1/3$ $\alpha=0$
Refer to caption Refer to caption Refer to caption
Figure 13. Distribution of the estimated CIL hyper-parameter $\hat{\theta}_{1}$ for the simulations described in Figure 1, considering strong ($\alpha=1$), weak $(\alpha=1/3)$ and no effect $(\alpha=0)$
$\alpha=1$ $\alpha=1/3$ $\alpha=0$
Refer to caption Refer to caption Refer to caption
Figure 14. Comparison of the CIL and ACPME prior inclusion log-odds ($\textup{logit}\pi_{j}({\bm{\theta}})$) for the simulations described in Figure 1, considering strong ($\alpha=1$), weak $(\alpha=1/3)$ and no effect $(\alpha=0)$. The y-axis shows the mean $\textup{logit}\pi_{j}({\bm{\theta}})$ per covariate type (confounders, instruments, predictors —active on the outcome but not on the treatment—, and spurious covariates), and the x-axis is the number of confounders (0 for no confounding, 6 for full confounding). Points in the figure are slightly offset on the $x$-axis when necessary to improve readability.
Proportion of confounders
$\alpha=1$ $\alpha=1/3$ $\alpha=0$
Refer to caption Refer to caption Refer to caption
Proportion of instruments
$\alpha=1$ $\alpha=1/3$ $\alpha=0$
Refer to caption Refer to caption Refer to caption
Figure 15. Proportion of confounders (top panels) and instruments (bottom panels) selected by each method in the simulation scenarios described in Figure 1, considering strong ($\alpha=1$), weak $(\alpha=1/3)$ and no effect $(\alpha=0)$. For Bayesian methods, we report the average marginal posterior inclusion probability. The x-axis quantifies the amount of confounding, measured by the number of covariates that are truly related to both the outcome and the treatment
@@ 15.6. Further results for the simulation study
In this section we expand upon the single treatment simulation results shown in Section 5.1. Figure 11 decomposes the mean squared errors of all methods shown in Figure 1 into the corresponding squared bias and variance. Standard high-dimensional methods like LASSO and BMA suffer from high bias and variance. In contrast, specialized treatment effect methods like BAC, ACPME and double LASSO show little bias but suffer from higher variance, particularly in low confounding scenarios.
Figure 12 assesses the sensitivity of BAC and ACPME to their respective hyperparameters. The performance of BAC is very sensitive to its tuning parameter. For BAC, setting the tuning parameter to $\omega=\infty$ means that, for any covariate found to be associated with the treatment, one forces its inclusion into the outcome model. $\omega=1$ means that inclusion in the treatment and outcome models is independent a priori, whereas $\omega=10$ represents a middle ground between the two other hyper-parameter choices. Regarding ACPME, its performance was fairly robust to its tuning parameter choice (related to the use of eigenvalues, correlations or a projection to measure associations).
Figure 13 shows the distribution of the CIL hyper-parameter $\hat{\theta}_{1}$ for the same simulation scenarios. Recall that $\theta_{1}>0$ is interpreted as high confounding, $\theta=0$ as neutral confounding, and $\theta_{1}<0$ as no confounding. As expected, regardless of the treatment effect size $\alpha$, CIL estimates $\hat{\theta}<0$ when there is no overlap between covariates that truly affect the outcome and those that truly affect the treatment (no confounding), and it estimates larger $\hat{\theta}_{1}$ as said overlap increases (up to full confounding, when all 6 truly active covariates in both equations overlap).
Figure 14 compares prior inclusion probabilities between CIL and ACPME, to illustrate their key distinction: the former adapts to the true amount of confounding (by using the outcome data) and the latter does not. ACPME favors equally the inclusion of confounders and of instruments, relative to predictors (covariates only associated to the outcome) and spurious covariates. Critically, this occurs to the same extent regardless of whether there truly is no confounding or high confounding. CIL, on the other hand, adapts to the true level of confounding. Under high confounding where there are more confounders than instruments (x-axis $\geq 5$ in Figure 14), inclusion of these two covariate types is encouraged. Under low confounding, where there are more instruments than confounders (x-axis $\leq 1$), their inclusion is disencouraged. Note also that CIL discouraged the inclusion of covariates unrelated to the treatment (whether or not related to the outcome), particularly in high confounding, this is because CIL also learns the overall level of sparsity via the intercept $\theta_{0}$.
Figure 15 (top panels) shows the proportion of confounders selected by each method. In high-confounding scenarios, BMA and to a lesser extent LASSO fail to include a fraction of the confounders, explaining the higher bias and variance observed in Figure 11, whereas the remaining methods include all counfounders. In settings with less or no confounding, all methods successfully included all confounders. The bottom panels show that double LASSO, BAC and ACPME included essentially all instruments, explaining their higher variance in Figure 11. LASSO also included a fair fraction of instruments, whereas CIL and BMA included less. BMA was particularly effective in this regard. Essentially, by inducing sparse solutions it includes less instruments, at the cost of missing some confounders.
Figure 16 summarizes model selection results for the simulations in Figure 1.
15.7. Simulations under growing dimensionality ($T=1$)
@@
15.8. Testing CIL to different amounts of confounders for $T=1$
@@
Figure 18 shows the effect of having various amounts of active confounders. The results look consistent to the effects reported in Figures 1 and 17, which are magnified for large amounts of active confounders. These are really challenging situations to tackle since the tested methods aim at model sparsity, while the true model size is relatively large. Although our method still performed at oracle rates in low-confounding scenarios, its relative performance is compromised for the highest levels of confounding. This occurred in part because accurate point estimation in (2.4) became increasingly harder as the correlation between covariates strengthened, which in turn influenced the ability of the algorithm to calibrate ${\bm{\theta}}$ reliably. Even in these hard cases, however, its performance is not excessively far to the best competing method, while it clearly outperformed BMA on all of them.